OpenStudy (anonymous):

If you buy a computer directly from the manufacturer for $2,273 and agree to repay it in 48 equal monthly installments at an annual interest rate of 21.12% on the unpaid balance, how much total interest will be paid on this amortized loan? (Round the amount of the installments to the nearest cent before calculating the final answer.)

OpenStudy (anonymous):

hmm okaii let me look over this ;)

OpenStudy (anonymous):

@some_someone do you know how to solve this?

jimthompson5910 (jim_thompson5910):

I think I have something, but I'm not 100% sure do you have answer choices?

OpenStudy (anonymous):

hmmm choices would be nice?

OpenStudy (anonymous):

no answer choices unfortunately

jimthompson5910 (jim_thompson5910):

well anyways, I found this formula and calculator that calculates the total cost of a loan (given initial amount borrowed, interest rate, and total life of loan) http://www.calcunation.com/calculators/business%20and%20finance/vehicle-total-cost.php The formula is Loan Total Cost = (c*L*n)/(1 - (1+c)^(-n)) Where L = loan principal amount (basically the initial balance you need to pay off) c = monthly interest rate (c = r/12) n = number of payment periods (ie number of months) In this case: L = 2273 c = 0.0176 (c = r/12 = 0.2112/12 = 0.0176) n = 48 (ie number of months) Note: n = 48 months = 4 years Loan Total Cost = (c*L*n)/(1 - (1+c)^(-n)) Loan Total Cost = (0.0176*2273*48)/(1 - (1+0.0176)^(-48)) Loan Total Cost = (0.0176*2273*48)/(1 - (1.0176)^(-48)) Loan Total Cost = (0.0176*2273*48)/(1 - 0.432811988064391) Loan Total Cost = (0.0176*2273*48)/(0.567188011935609) Loan Total Cost = (1920.2304)/(0.567188011935609) Loan Total Cost = 3385.52712608813 Loan Total Cost = 3385.53 ------------------------------------------------------- Now use this to find the interest Interest = Total Cost - Initial Balance Interest = 3385.53 - 2273 Interest = 1112.53

jimthompson5910 (jim_thompson5910):

not 100% sure because I've used that formula before and it's been off by a few cents before, so idk, it's hard to say

OpenStudy (anonymous):

ill see if it's correct soon, thanks for the help!

jimthompson5910 (jim_thompson5910):

you're welcome

jimthompson5910 (jim_thompson5910):

and thx for letting me know